# Mitigating Label Noise using Prompt-Based Hyperbolic Meta-Learning in Open-Set Domain Generalization

### v1

@@ 4 Method
Meta-learning proves to be highly effective domain generalization by simulating a diverse range of cross-domain tasks during the training process, thereby facilitating the learning of adaptable representations capable of generalizing to previously unseen domains [32]. By emphasizing task-specific and category-specific learning dynamics, meta-learning enables models to efficiently adapt to novel domain and novel class distributions and effectively address the challenges posed by domain shifts under open-set conditions [66, 57]. In this work we improve the robustness of meta-learning for open-set domain generalization against label noise perturbation during the training time. In this section, we will introduce our proposed novel method, HyProMeta, which is composed of hyperbolic categorical prototype-based label noise-aware meta-learning (as introduced in Section 4.1) and new category-aware prompt learning to improve the discriminative ability of the learned embeddings (as introduced in Section 4.2).
@@
1:$\Omega_{\mathcal{D}_{S}}\coloneqq\text{Set of source domains}$,  $\mathcal{C}\coloneqq\text{Set of known classes},\quad C\coloneqq|\mathcal{C}|$,
@@
2:$k\in\{1,\ldots,C\}\coloneqq\text{Class index}$,  $\mathbf{M}_{\alpha}\coloneqq\text{Neural network backbone with parameters }\alpha$,
3:$\mathbf{H}_{\beta}\coloneqq\text{Classification head with parameters $\beta$}$,
[…]
5:
6:while not converged do
7:     $\{\mathcal{D}_{s_{i}},\mathcal{D}_{s_{j}}\}\leftarrow\Omega_{\mathcal{D}_{S}}$triangleright Select two random source domains
[…]
@@
9:     $\{\mathbf{P}_{k}^{s},\hat{d}_{k}^{s}\}\leftarrow\text{HyperbolicPrototypesAndThresholds}(\{\mathcal{D}_{s_{i}},\mathcal{D}_{s_{j}}\})$triangleright Domain- and category-specific
@@
10:     $d_{k}^{s}\leftarrow\text{ComputeHyperbolicDistances}(\{\mathcal{D}_{s_{i}},\mathcal{D}_{s_{j}}\},\mathbf{P}_{k}^{s})$triangleright Calculate distances to prototypes
11:     $\Omega_{c},\Omega_{n}\leftarrow\{x|d_{k}^{s}(x)<\hat{d}_{k}^{s}\},\{x|d_{k}^{s}(x)\geq\hat{d}_{k}^{s}\}$triangleright Split into clean and noisy set.
@@
12:     $\mathbf{y}_{n}^{*}\leftarrow\operatorname*{arg\,min}_{k}d_{k}^{s_{i}}(\Omega_{n})$triangleright Correct noisy labels using nearest prototype
13:     $\Omega_{n}^{*}\leftarrow\{\Omega_{n},\mathbf{y}_{n}^{*}\}$triangleright Create corrected noisy set
[…]
@@
15:     $\{\mathbf{B}^{s_{i}}_{n},\mathbf{B}^{s_{i}}_{c}\}\leftarrow\text{SampleBatches}(\Omega_{n}^{*},\Omega_{c}^{s_{i}})$triangleright Sample from noisy and clean sets
@@
16:     $\hat{\mathbf{B}}^{s_{i}}_{c}\leftarrow\text{SampleDifferentClasses}(\Omega_{c}^{s_{i}},\text{Classes}(\mathbf{B}^{s_{i}}_{c}))$triangleright Sample batch with different classes
[…]
18:     $\mathbf{B}_{mixed}\leftarrow\text{Mix}(\mathbf{B}^{s_{i}}_{c},\hat{\mathbf{B}}^{s_{i}}_{c})$triangleright Mix clean samples
@@
19:     $\hat{\mathbf{B}}^{s_{i}}_{a}\leftarrow\text{PromptAug}(\mathbf{B}_{mixed},\mathbf{p}_{l},\mathbf{M}_{crop})$triangleright Apply prompt augmentation
@@
20:     $\mathbf{y}_{a}\leftarrow(C+1)\cdot\mathbf{1}$triangleright Assign unknown class label to augmented samples
@@
21:     $L_{m-train}\leftarrow CE(\mathbf{H}_{\beta}(\mathbf{M}_{\alpha}(\mathbf{B}^{s_{i}}_{c})),\mathbf{y}_{c})$triangleright Clean samples loss
@@
22:     $L_{m-train}\leftarrow L_{m-train}+CE(\mathbf{H}_{\beta}(\mathbf{M}_{\alpha}(\hat{\mathbf{B}}^{s_{i}}_{a})),\mathbf{y}_{a})$triangleright Augmented samples loss
@@
23:     $L_{m-train}\leftarrow L_{m-train}+CE(\mathbf{H}_{\beta}(\mathbf{M}_{\alpha}(\mathbf{B}^{s_{i}}_{n})),\mathbf{y}_{n}^{*})$triangleright Corrected noisy samples loss
[…]
@@
26:     $\mathbf{B}_{c}^{s_{i}}\leftarrow\text{SampleBatch}(\Omega_{c}^{s_{i}})$triangleright Sample from first domain
@@
27:     $\mathbf{B}_{c}^{s_{j}}\leftarrow\text{SampleBatch}(\Omega_{c}^{s_{j}})$triangleright Sample from second domain
@@
28:     $L_{m-test}\leftarrow CE(\mathbf{H}_{\beta}(\mathbf{M}_{\alpha}(\mathbf{B}_{c}^{s_{i}})),\mathbf{y}_{c}^{s_{i}})$triangleright First domain loss
@@
29:     $L_{m-test}\leftarrow L_{m-test}+CE(\mathbf{H}_{\beta}(\mathbf{M}_{\alpha}(\mathbf{B}_{c}^{s_{j}})),\mathbf{y}_{c}^{s_{j}})$triangleright Second domain loss
30:
[…]
32:end while
@@ 4.1 Hyperbolic Categorical Prototype-Based Meta Learning
Hyperbolic Prototype Calculation: Hyperbolic space enhances the generalization capabilities of deep learning models by efficiently modeling hierarchical information and enabling compact representations of complex relationships. Its intrinsic geometric properties facilitate the preservation of both global and local data structures, reducing overfitting and improving robustness to unseen data [14, 44].
Leveraging these generalizability advantages, we utilize hyperbolic space to generate categorical prototypes from the extracted embeddings of training samples during meta-learning.
A hyperbolic space is defined as a complete and connected mutually isometric Riemannian manifold, which is characterized by a constant negative sectional curvature [17, 1]. The Poincaré ball model is highly effective in hyperbolic space modeling, preserving both global and local geometric relationships [50]. We select the Poincaré ball model which is introduced in the Eq. 1.
@@
where $\gamma=\frac{1}{r^{2}}$. The mapping from Euclidean space to hyperbolic space can be represented by Eq. 4 and the base point $\mathbf{q}$ is set as zeros. $\lambda_{z}$ denotes a scale factor.
@@
Given input $\mathbf{x}$, we first pass it through the leveraged deep learning backbone $\mathbf{M}_{\alpha}$ to obtain high dimensional embeddings $\mathbf{z}$, where $\mathbf{z}=\mathbf{M}_{\alpha}(\mathbf{x})$. Then we project the resultant embedding into hyperbolic space using $\hat{\mathbf{z}}=\Phi_{h}(\mathbf{z})$. We first calculate the hyperbolic categorical center according to the following equation for each domain $s$ and category $k$ ($s\in\Omega_{s}$ and $k\in\Omega_{k}$).
@@
where $\mathbf{e}_{k}$ denotes the hyperbolic center of category $k$. $N_{k}$ denotes the number of samples inside category $k$. The hyperbolic categorical center is calculated with a fixed epoch step $N_{epoch}$. After the calculation of the domain-agnostic categorical prototypes in the hyperbolic space, we achieve different data partitions for meta-test and meta-train assignment using the hyperbolic distance of each sample to its corresponding domain-agnostic categorical prototype. The distance can be represented by Eq. 6.
@@
where $i$ indicates the sample index, $k$ denotes the category, and $s$ denotes one source domain. Then, we calculate the mode of distances within each domain for each category, denoted by $\hat{d}_{k}^{s}$, where $\hat{d}_{k}^{s}=\text{Mode}(\{d_{k}^{(i,s)}\})$. The samples that have smaller hyperbolic distances to their corresponding domain-agnostic categorical prototypes are assigned to the clean set while the others are assigned to the noisy set, where we utilize the hyperbolic distances as the measurement of label uncertainty. The samples from the noisy set are relabeled by the category of its nearest hyperbolic prototype. Then we use the aforementioned partitions as described in Alg. 1 to achieve label noise agnostic meta-learning to handle the OSDG-NL challenge.
@@
To further enhance generalization capabilities, the method integrates a learnable prompt for category-aware data augmentation as an additional category, effectively capturing intra-class variability and reinforcing the model’s robustness to distribution shifts, which will introduced in detail in the next subsection.
@@ 4.2 New Category Aware Prompt Learning (NCA-Prompt)
We have previously introduced the hyperbolic prototype-based meta-learning framework, which separates meta-train and meta-test phases by dynamically partitioning data into clean and noisy subsets based on their respective hyperbolic distances from categorical prototypes. While this framework effectively manages label shifts induced by label noise, it is also necessary to address distribution shifts that occur within a single class and domain. Specifically, the model must demonstrate high confidence when encountering OOD data that remain within the same category and domain.
Existing approaches, such as Deep Metric Learning (DML) [63], focus on achieving compact intra-class representations. However, DML methods are inherently sensitive to label noise. Forcing samples with noisy or incorrect labels to align closely in the feature space can degrade the performance of our hyperbolic prototype-based meta-learning framework, as it inadvertently propagates noise and hinders the model’s ability to distinguish between clean and noisy samples. Therefore, an alternative strategy is required to achieve robust intra-class compactness during the meta-train process while being resilient to noisy labels.
Enhancing the model’s generalizability to OOD samples beyond mere label shifts is equally critical. While MixStyle [83] has shown that domain-mixed samples can improve generalization across different domains via data augmentation, it does not explicitly target the intra-class and intra-domain generalizability required in our scenario. To bridge this gap, we propose a novel categorical-aware prompt learning mechanism to complement hyperbolic prototype-based meta-learning, particularly in the context of open-set domain generalization under noisy labels.
@@
Additionally, we introduce learnable prompt $\mathbf{p}_{l}\in\mathcal{R}^{H\times W\times 3}$, which has the same spatial dimensions as the input image. The learnable prompt is used to augment the mixed samples in order to enrich the data distribution for different categories during the training phase. A random cropping window is selected on the mixed sample $\mathbf{z}^{*}$, and the corresponding region in $\mathbf{z}^{*}$ is replaced with the same region from $\mathbf{p}_{l}$. This operation is formally expressed as Eq. 7.
@@ 5 Experiments
In this section, we first illustrate the implementation details of our approach in Section 5.1, then we deliver the analysis of the constructed benchmarks and our proposed method in Section 5.2, followed by the analysis of ablation studies (in Section 5.3), analysis of the t-SNE visualizations (in Section 5.4), and analysis of the confidence score towards seen and unseen categories in the test domain (in Section 5.5). The ablation of NCA-Prompt, HYB-Meta, and hyperparameter $N_{epoch}$ are analyzed in Section 5.6, Section 5.7, and Section 5.8.
@@ 5.1 Implementation Details
All experiments are conducted using PyTorch 2.0 and a single NVIDIA A100 GPU. Training is capped at $1\times 10^{4}$ steps, employing the SGD optimizer with a Learning Rate (LR) of $1\times 10^{-3}$ and a batch size of $16$. A learning rate decay of $1\times 10^{-1}$ is applied after $8\times 10^{3}$ meta-training steps. The experiments use a worker number of $4$, and the value of $\gamma$ is fixed at $2\times 10^{-5}$. For feature extraction, the ConvNet [84] is employed as the backbone network on the DigitsDG dataset, following the setup outlined in [84]. Hyperparameter $N_{epoch}$ is set as $500$, which is chosen according to the ablation. OSCR is chosen as major evaluation metric while H-score and close-set accuracy (ACC) are chosen as secondary metrics following [66].
Table 1: Results (%) of PACS on ResNet18 [24]. The open-set ratio is $6{:}1$ and symmetric label noise with ratio $20\%$ is selected.

### v2

@@ 4 Method
Meta-learning proves to be highly effective for domain generalization by simulating a diverse range of cross-domain tasks during the training process, thereby facilitating the learning of adaptable representations capable of generalizing to previously unseen domains [li2018learning]. By emphasizing task-specific and category-specific learning dynamics, meta-learning enables models to efficiently adapt to novel domains and novel class distributions and effectively address the challenges posed by domain shifts under open-set conditions [wang2023generalizable, shu2021open]. In this work we improve the robustness of meta-learning for open-set domain generalization against label noise perturbation during the training time. In this section, we will introduce our proposed novel method, HyProMeta, which is composed of hyperbolic categorical prototype-based label noise-aware meta-learning (as introduced in Section 4.1) and new category-aware prompt learning to improve the discriminative ability of the learned embeddings (as introduced in Section 4.2).
@@
1: $\Omega_{\mathcal{D}_{S}}\coloneqq\text{Set of source domains}$,  $\mathcal{C}\coloneqq\text{Set of known classes},\quad C\coloneqq|\mathcal{C}|$, $k\in\{1,\ldots,C\}\coloneqq\text{Class index}$,  $\mathbf{M}_{\alpha}\coloneqq\text{Neural network backbone with parameters }\alpha$, $\mathbf{H}_{\beta}\coloneqq\text{Classification head with parameters $\beta$}$, $CE\coloneqq\text{Cross entropy loss function}$,  $\mathbf{p}_{l}\coloneqq\text{Learnable image-wise prompt}$
2: while not converged do
3:   $\{\mathcal{D}_{s_{i}},\mathcal{D}_{s_{j}}\}\leftarrow\Omega_{\mathcal{D}_{S}}$ //˜Select two random source domains 4:
5:   $\triangleright$ Hyperbolic Distance Based Label Correction   $\triangleleft$
@@
6:   $\{\mathbf{P}_{k}^{s},\hat{d}_{k}^{s}\}\leftarrow\text{HyperbolicPrototypesAndThresholds}(\{\mathcal{D}_{s_{i}},\mathcal{D}_{s_{j}}\})$ //˜Domain- and category-specific
@@
7:   $d_{k}^{s}\leftarrow\text{ComputeHyperbolicDistances}(\{\mathcal{D}_{s_{i}},\mathcal{D}_{s_{j}}\},\mathbf{P}_{k}^{s})$ //˜Calculate distances to prototypes
8:   $\Omega_{c},\Omega_{n}\leftarrow\{x|d_{k}^{s}(x)<\hat{d}_{k}^{s}\},\{x|d_{k}^{s}(x)\geq\hat{d}_{k}^{s}\}$ //˜Split into clean and noisy set.
@@
9:   $\mathbf{y}_{n}^{*}\leftarrow\operatorname*{arg\,min}_{k}d_{k}^{s_{i}}(\Omega_{n})$ //˜Correct noisy labels using nearest prototype
10:
$\Omega_{n}^{*}\leftarrow\{\Omega_{n},\mathbf{y}_{n}^{*}\}$
//˜Create corrected noisy set
11:
12:   $\triangleright$ Meta-train on first domain   $\triangleleft$
@@
13:   $\{\mathbf{B}^{s_{i}}_{n},\mathbf{B}^{s_{i}}_{c}\}\leftarrow\text{SampleBatches}(\Omega_{n}^{*},\Omega_{c}^{s_{i}})$ //˜Sample from noisy and clean sets
@@
14:   $\hat{\mathbf{B}}^{s_{i}}_{c}\leftarrow\text{SampleDifferentClasses}(\Omega_{c}^{s_{i}},\text{Classes}(\mathbf{B}^{s_{i}}_{c}))$ //˜Sample batch with different classes
15:   $\mathbf{M}_{crop}\leftarrow\text{GenerateRandomMask}()$ //˜Generate random crop mask
16:   $\mathbf{B}_{mixed}\leftarrow\text{Mix}(\mathbf{B}^{s_{i}}_{c},\hat{\mathbf{B}}^{s_{i}}_{c})$ //˜Mix clean samples
@@
17:   $\hat{\mathbf{B}}^{s_{i}}_{a}\leftarrow\text{PromptAug}(\mathbf{B}_{mixed},\mathbf{p}_{l},\mathbf{M}_{crop})$ //˜Apply prompt augmentation
@@
18:   $\mathbf{y}_{a}\leftarrow(C+1)\cdot\mathbf{1}$ //˜Assign unknown class label to augmented samples
@@
19:   $L_{m-train}\leftarrow CE(\mathbf{H}_{\beta}(\mathbf{M}_{\alpha}(\mathbf{B}^{s_{i}}_{c})),\mathbf{y}_{c})$ //˜Clean samples loss
@@
20:   $L_{m-train}\leftarrow L_{m-train}+CE(\mathbf{H}_{\beta}(\mathbf{M}_{\alpha}(\hat{\mathbf{B}}^{s_{i}}_{a})),\mathbf{y}_{a})$ //˜Augmented samples loss
@@
21:   $L_{m-train}\leftarrow L_{m-train}+CE(\mathbf{H}_{\beta}(\mathbf{M}_{\alpha}(\mathbf{B}^{s_{i}}_{n})),\mathbf{y}_{n}^{*})$ //˜Corrected noisy samples loss
22:   $\text{UpdateParameters}(L_{m-train})$ //˜Update based on meta-train loss 23:
24:   $\triangleright$ Meta-test on both domains   $\triangleleft$
@@
25:   $\mathbf{B}_{c}^{s_{i}}\leftarrow\text{SampleBatch}(\Omega_{c}^{s_{i}})$ //˜Sample from first domain
@@
26:   $\mathbf{B}_{c}^{s_{j}}\leftarrow\text{SampleBatch}(\Omega_{c}^{s_{j}})$ //˜Sample from second domain
@@
27:   $L_{m-test}\leftarrow CE(\mathbf{H}_{\beta}(\mathbf{M}_{\alpha}(\mathbf{B}_{c}^{s_{i}})),\mathbf{y}_{c}^{s_{i}})$ //˜First domain loss
@@
28:   $L_{m-test}\leftarrow L_{m-test}+CE(\mathbf{H}_{\beta}(\mathbf{M}_{\alpha}(\mathbf{B}_{c}^{s_{j}})),\mathbf{y}_{c}^{s_{j}})$ //˜Second domain loss 29:
30:   $\text{UpdateParameters}(L_{m-test}+L_{m-train})$ //˜Final parameter update
@@ 4.1 Hyperbolic Categorical Prototype-Based Meta Learning
Hyperbolic Prototype Calculation: Hyperbolic space enhances the generalization capabilities of deep learning models by efficiently modeling hierarchical information and enabling compact representations of complex relationships. Its intrinsic geometric properties facilitate the preservation of both global and local data structures, reducing overfitting and improving robustness to unseen data [cui2024rethinking_few_shot, liu2020hyperbolic].
Leveraging these generalizability advantages, we utilize hyperbolic space to generate category prototypes from the extracted embeddings of training samples during meta-learning.
A hyperbolic space is defined as a complete and connected mutually isometric Riemannian manifold, which is characterized by a constant negative sectional curvature [fang2023Poincare, absil2008optimization]. The Poincaré ball model is highly effective in hyperbolic space modeling, preserving both global and local geometric relationships [mishne2023numerical]. We select the Poincaré ball model, which is introduced in Eq. 1.
@@
where $\gamma=\frac{1}{r^{2}}$. The mapping from Euclidean space to hyperbolic space can be represented by Eq. 4, and the base point $\mathbf{q}$ is set to zeros. $\lambda_{z}$ denotes a scale factor.
@@
Given input $\mathbf{x}$, we first pass it through the leveraged deep learning backbone $\mathbf{M}_{\alpha}$ to obtain high dimensional embeddings $\mathbf{z}$, where $\mathbf{z}=\mathbf{M}_{\alpha}(\mathbf{x})$. Then we project the resultant embedding into hyperbolic space using $\hat{\mathbf{z}}=\Phi_{h}(\mathbf{z})$. We calculate the hyperbolic category center according to Eq. 5 for each domain $s$ and category $k$, where $s\in\Omega_{s}$ and $k\in\Omega_{k}$.
@@
where $\mathbf{e}_{k}$ denotes the hyperbolic center of category $k$. $N_{k}$ denotes the number of samples inside category $k$. The hyperbolic category center is calculated with a fixed epoch step $N_{epoch}$. After the calculation of the domain-agnostic category prototypes in the hyperbolic space, we achieve different data partitions for meta-test and meta-train assignment using the hyperbolic distance of each sample to its corresponding domain-agnostic category prototype. The distance can be represented by Eq. 6.
@@
where $i$ indicates the sample index, $k$ denotes the category, and $s$ denotes one source domain. Then, we calculate the mode of distances within each domain for each category, denoted by $\hat{d}_{k}^{s}$, where $\hat{d}_{k}^{s}=\text{Mode}(\{d_{k}^{(i,s)}\})$, where the distances are binned to their upper bound using binned resolution of $\delta$. We use the mode of hyperbolic distances for clean/noisy partitioning because it effectively captures the dominant distribution of clean samples within each class and domain, which typically forms a compact cluster. Unlike mean or median, the mode is less affected by noisy outliers and takes the majority structure of clean-label samples into consideration, which enables more robust thresholding under severe label noise. The samples that have smaller hyperbolic distances to their corresponding domain-agnostic category prototypes are assigned to the clean set, while the others are assigned to the noisy set, where we utilize the hyperbolic distances as the measurement of label uncertainty. The samples from the noisy set are relabeled by the category of their nearest hyperbolic prototype. Then we use the aforementioned partitions as described in Alg. 1 to achieve label noise agnostic meta-learning to handle the OSDG-NL challenge.
@@
To further enhance generalization capabilities, the method integrates a learnable prompt for category-aware data augmentation as an additional category, effectively capturing intra-class variability and reinforcing the model’s robustness to distribution shifts, which will be introduced in detail in the next subsection.
@@ 4.2 New Category Aware Prompt Learning (NCA-Prompt)
We have previously introduced the hyperbolic prototype-based meta-learning framework, which separates meta-train and meta-test phases by dynamically partitioning data into clean and noisy subsets based on their respective hyperbolic distances from category prototypes. While this framework effectively manages label shifts induced by label noise, it is also necessary to address distribution shifts that occur within a single class and domain. Specifically, the model must demonstrate high confidence when encountering OOD data that remain within the same category and domain.
Existing approaches, such as Deep Metric Learning (DML) [wang2017deep], focus on achieving compact intra-class representations. However, DML methods are inherently sensitive to label noise. Forcing samples with noisy or incorrect labels to align closely in the feature space can degrade the performance of our hyperbolic prototype-based meta-learning framework, as it inadvertently propagates noise and hinders the model’s ability to distinguish between clean and noisy samples. Therefore, an alternative strategy is required to achieve robust intra-class compactness during the meta-train process while being resilient to noisy labels.
Enhancing the model’s generalizability to OOD samples beyond mere label shifts is equally critical. While MixStyle [zhou2020domain] has shown that domain-mixed samples can improve generalization across different domains via data augmentation, it does not explicitly target the intra-class and intra-domain generalizability required in our scenario. To bridge this gap, we propose a novel category-aware prompt learning mechanism to complement hyperbolic prototype-based meta-learning, particularly in the context of open-set domain generalization under noisy labels.
@@
Additionally, we introduce a learnable prompt $\mathbf{p}_{l}\in\mathcal{R}^{H\times W\times 3}$, which has the same spatial dimensions as the input image. The learnable prompt is used to augment the mixed samples in order to enrich the data distribution for different categories during the training phase. A random cropping window is selected on the mixed sample $\mathbf{z}^{*}$, and the corresponding region in $\mathbf{z}^{*}$ is replaced with the same region from $\mathbf{p}_{l}$. This operation is formally expressed as Eq. 7.
@@
To generate a rectangle prompt region, two random points within the image dimensions are sampled to define the top-left and bottom-right corners. The rectangle’s area is calculated and compared against the total image area to ensure that its relative size falls within the specified range $\left[rel_{min},rel_{max}\right]$. This ensures that the selected region for each object occupies a controllable proportion of the image, maintaining both variability and constraint in the prompt generation.
@@ 5 Experiments
In this section, we first illustrate the implementation details of our approach in Section 5.1 and evaluation metrics in Section 5.2. Then we deliver the analysis on the constructed benchmarks and our proposed method in Section 5.3. Next, we conduct analysis towards the effect brought by different pretraining strategies in Section 5.4 and the analysis on the label-clean/noisy set partition in Section 5.5, followed by the analysis on the effect brought by different textual encoder for asymmetric label noise generation, ablation studies of different modules (in Section 5.7), analysis on the t-SNE visualizations (in Section 5.8), and analysis on the confidence score towards seen and unseen categories in the test domain (in Section 5.9). The ablation of NCA-Prompt, HYB-Meta, and hyperparameter $N_{epoch}$ are analyzed in Section 5.10, Section 5.11, and Section 5.12.
@@ 5.1 Implementation Details
All experiments are conducted using PyTorch 2.0 and a single NVIDIA A100 GPU. Training is capped at $1\times 10^{4}$ steps, employing the SGD optimizer with a Learning Rate (LR) of $1\times 10^{-3}$ and a batch size of $16$. A learning rate decay of $1\times 10^{-1}$ is applied after $8\times 10^{3}$ meta-training steps. $\delta$ is chosen as $1\times 10^{-2}$. The experiments use a worker number of $4$, and the value of $\gamma$ is fixed at $2\times 10^{-5}$. For feature extraction, the ConvNet [zhou2021domain] is employed as the backbone network on the DigitsDG dataset, following the setup outlined in [zhou2021domain]. $\left[rel_{min},rel_{max}\right]$ is set as $\left[0.2,0.6\right]$. The curvature, $\lambda_{z}$, and radius of the Poincaré ball are selected as $5\times 10^{-2}$, $2.0$, and $9.99$, the dimension of $z$ is $512$. Hyperparameter $N_{epoch}$ is set as $500$, which is chosen according to the ablation. OSCR is chosen as the major evaluation metric, while H-score and close-set accuracy (ACC) are chosen as secondary metrics following [wang2023generalizable].
5.2 Evaluation Metrics
The Acc metric denotes the closed-set accuracy evaluated on seen categories and serves to measure the correctness of classification within the known label space. In contrast, the H-score and OSCR are widely adopted in the OSDG field [wang2022generalizing, peng2024advancing] to evaluate open-set recognition performance. Since the H-score requires a manually selected threshold derived from source domain validation sets to distinguish between seen and unseen categories, it is treated as a secondary metric in our evaluation. Instead, we adopt OSCR, initially proposed by MEDIC [wang2023generalizable], as our primary evaluation metric, as it does not depend on a fixed threshold and provides a more flexible and robust assessment of open-set recognition. To compute the H-score, a threshold ratio $\sigma$ is first selected to distinguish samples from seen and unseen categories. If the predicted confidence score for a sample falls below $\sigma$, it is classified as an unseen category. The accuracy for samples predicted as seen categories is then computed based on their ground-truth seen labels and denoted as $Acc_{k}$. For unseen category evaluation, a binary classification approach is applied: samples from seen categories are labeled as $1$ and those from unseen categories as 0, and the classification accuracy is denoted as $Acc_{u}$. The final H-score is calculated as:
$H_{score}=\frac{2Acc_{u}Acc_{k}}{Acc_{u}+Acc_{k}}.$ (8)
The OSCR metric evaluates the quality of confidence scores through a moving threshold mechanism, combining accuracy with AUROC-like behavior. Unlike AUROC, however, OSCR only considers correctly predicted samples at each threshold, blending the evaluation principles of both H-score and AUROC to better reflect model performance in the OSDG setting.
Table 1: Results (%) of PACS on ResNet18 [he2016deep]. The open-set ratio is 6:16{:}1 and symmetric label noise with a ratio of $20\%$ is selected.